# Thread: A Logic-Based Data Organization Paradigm for How-To Question Answering with Retrieval Augmented Generation

### v1

@@ 4.1 Scenarios and Datasets
We evaluate Thread on two open-domain scenarios: Web Navigation (Mind2Web Deng et al. (2023)) and Wikipedia Instructions (WikiHow Koupaee and Wang (2018)), and one industrial scenario: Incident Mitigation (IcM Shetty et al. (2022); An et al. (2024)). Appendix B provides one example of each dataset to better demonstrate the linear and dynamic how-to question. More details about datasets can be found in Appendix DEF.
Web Navigation. Mind2Web Deng et al. (2023) is a dataset designed for web agents to perform complex tasks on real-world websites based on language instructions. Each task is a dynamic how-to question, with multiple possible outcomes depending on the state of the executed actions.
Wikipedia Instructions. WikiHow [^4] https://www.wikihow.com is a platform containing numerous articles that offer step-by-step guidance on various procedural tasks, ranging from entertainment to technology. Each article is typically titled “How to” and includes a brief task description, followed by detailed steps. We use this dataset to demonstrate Thread’s performance on linear how-to questions.
Incident Mitigation. Incident Mitigation Shetty et al. (2022); An et al. (2024); Jiang et al. (2024) is essential for operating large-scale cloud services, where engineers use Troubleshooting Guides (TSGs) to address incidents. Each step in incident mitigation can yield different outcomes depending on the state, making it suitable for testing Thread on dynamic how-to questions. Appendix D provides details. Unlike the other open-domain datasets, we perform a human evaluation involving twenty on-call engineers (OCEs) [^5] Varying across new-hire and experienced OCEs. responsible for incident mitigation. We collect five incidents [^6] Mitigating one incident costs each OCE around 30 minutes to 1 hour manually, and we collect five incidents to ensure consistency, engagement, and ethical treatment., classified into two simple and three hard ones based on their mitigation steps in history. Each OCE is tasked with mitigating all five incidents, using randomly ordered baselines per incident to avoid familiarity bias. The QA system initiates automated mitigation for each incident; if it encounters a failure at any step, an OCE intervenes to address the issue before the system resumes automated procedures.
@@ 4.2 Documents for Retrieval
Since the Mind2Web datasets lack relevant documents, we created documents for retrieval to fit the RAG-based QA system. Assuming the same website has help docs applicable to all tasks in the website, we utilized the “Cross-Task” test set and selected examples from the training set to create informative documents for each website, following Wang et al. (2023). This allows the generation of documents for retrieval to effectively accomplish tasks within the RAG-based QA system.
For the WikiHow dataset, we used publicly available Windows Office Support Docs [^7] https://github.com/MicrosoftDocs/OfficeDocs-Support as retrieval documents. We selected around 100 tasks from the WikiHow dataset tagged with Microsoft products like Word, PowerPoint, and Teams, each containing 10 to 40 steps related to Windows operations.
For the IcM dataset, we collected 56 TSGs as documents from an enterprise-level engineering team responsible for a large-scale cloud platform. The selected incidents in the IcM dataset can be resolved using knowledge from these TSGs.
Table 1 shows the dataset statistics. Note that in IcM dataset, one OCE’s data was contaminated during the experiment, so we removed that OCE’s data. It results in 19 OCEs with 95 tasks.
@@
Table 1: The statistics of datasets, including the number of documents, tasks, steps, and derived chunks and LUs.
@@ 4.3 Baselines
In the Mind2Web dataset, previous work has not treated it as how-to questions. State-of-the-art methods like SYNASE Zheng et al. (2023) and MINDACT Deng et al. (2023) either use In-Context Learning (ICL), providing few-shot demonstrations, or Supervised Learning (SL) to finetune a model on the training set. To ensure a fair comparison with our LLM endpoints [^8] We use GPT-3.5 and GPT-4 with version 1106-preview., we re-implemented the MINDACT experiments with the same demonstrations and included chat history as extra context. In our paper, we treat the Mind2Web dataset as dynamic how-to questions and use our RAG-based system to solve these tasks. For comparison, we use doc-based (providing the entire document) and chunk-based data organization paradigms as baselines against Thread. This RAG-based QA system with different data organization paradigms also serves as baselines for the WikiHow and IcM datasets.
@@ 4.4 Evaluation Metrics
For the Mind2Web dataset, each task is treated as a multi-choice question. We adopt the evaluation metrics used in Deng et al. (2023), which include: Element Accuracy (Ele. Acc) to compare the chosen HTML element with all provided elements; Operation F1 (Op. F1) to calculate the token-level F1 score for predicted operations such as “CLICK”, “TYPE IN”, etc.; Step Success Rate (Step SR), where a step is successful if both the selected element and predicted operation are accurate; and Success Rate (SR), where a task is successful only if all steps are successful.
For the WikiHow dataset, which contains ground truth steps, we leverage LLMs to extract “Action Items” from each ground truth step and generated step, and we use the following metrics: Precision ($P=\frac{\#matched\_items}{\#total\_generated\_items}$); Recall ($R=\frac{\#matched\_items}{\#total\_groundtruth\_items}$); F1; and Success Rate (SR) to assess if the generated steps can successfully complete the task, using LLMs to evaluate (Appendix F shows the evaluation prompt).
@@ 5 Experimental Results
Web Navigation. The overall performance on Mind2Web is shown in Table 2, where we compare our method with baselines and both doc-based and chunk-based RAG methods. We observe that providing informative documents (regardless of the data organization paradigm) within RAG methods significantly improves performance. RAG methods outperform ICL and SL methods. By incorporating external documents, the doc-based RAG method achieves performance comparable to the best results of MINDACT. Notably, our Thread paradigm is the best among RAG methods, showing improvements of 4.06% in Ele. Acc, 3.49% in Step SR and 3.57% in SR. Note that MINDACT-SL gets the highest Op. F1 due to label distribution imbalance, leading the model to favor generating the most frequent operations.
@@
Table 2: Experiment results on Mind2Web. “*” represents taking results from the original paper.
Wikipedia Instructions. Table 3 presents the performance of various data organization paradigms on WikiHow tasks, experimenting with both single-turn and multi-turn interactions. In the single-turn setting, where the QA system generates the entire plan in one interaction, our Thread-based method outperforms the doc-based method which provides the entire relevant document, indicating that Thread offers information with less redundancy. The chunk-based method is not included as it failed to retrieve enough relevant chunks using the question as the query.
In the multi-turn setting, the QA system performs iterative retrieval, which shows superior performance compared to single-turn. This indicates the advantage of a step-by-step approach in handling how-to questions. Notably, the SR improves significantly from 58.76% to 68.04% in doc-based paradigm and from 63.91% to 72.16% in Thread paradigm. Within multi-turn setting, chunk-based method divides documents into chunks which disrupts internal connections, leading to a low SR of 20.62%. In contrast, both doc-based and Thread-based paradigms perform significantly better. Our Thread paradigm excels across all metrics, achieving the highest SR of 72.16%, highlighting its effectiveness in maintaining and modeling the connections between steps in linear how-to questions.
@@
Table 3: Experiment results on WikiHow with different interaction manners and data organization paradigms.
Incident Mitigation. Figure 4 illustrates the advantages of Thread over other paradigms when addressing complex dynamic how-to questions. Both chunk-based and doc-based RAG methods exhibit comparable ineffectiveness in incident mitigation. In contrast, our Thread paradigm achieves the best score across all metrics, particularly achieving a significant increase in SR from 21.02% to 33.33%. Furthermore, P.F. Step SR indicates our method’s capability to handle significantly more steps before encountering a failure, thus reducing the need for human intervention. Consequently, our method achieves the highest Step SR and requires the fewest interaction Turns to mitigate both simple and complex incidents.
Figure 4: Experiment results on simple and hard incidents of IcM dataset with OCEs.
5.1 Ablation on QA System Settings
This section presents an ablation study on the key settings of a RAG-based QA system. While retriever and generator variants have been explored in other research Gao et al. (2023), we use text-embedding-ada-002 OpenAI (2022) for the retriever and GPT-4 as the LLM-based generator. The ablation study is performed in Mind2Web dataset.
Multi-turn Interaction. As shown in Table 3, the multi-turn setting is better than single-turn setting in answering how-to questions. We use multi-turn setting for the other scenarios in this paper.
Chat History. The chat history helps the system to check previous actions and results, in order to help the system make better decisions. Without chat history, the performance drops as shown in Appendix E.3. We include chat history for all RAG-based methods in the experiments.
Retrieval Units Selector. As aforementioned in §3.4 and Figure 3, the Selector picks the most relevant retrieval units from the top-K retrieved units. Table 4 shows the ablation of the retrieval unit selector. We compare chunk and LU as retrieval units, finding that chunk selection degrades the performance of chunk-based RAG, reducing Ele. Acc by 4.29% and Op.F1 by 3.38%. Conversely, the selector improves all metrics in the Thread-based RAG method. Unlike LU selection, which filters out irrelevant LUs with prerequisites, chunk selection does not consider connections between chunks and may filter out relevant chunks. Therefore, we enable the selector only for Thread-based method.
@@ 5.1 Ablation on QA System Settings
Table 4: Ablation study of retrieval unit selection.
Multi-Agent System. We investigate the multi-agent design in QA system. As detailed in Appendix E.4, the multi-agent RAG system Wu et al. (2023); An et al. (2024) outperforms single-agent systems, where only the LLM-based generator acts as the agent, in resolving how-to questions.
@@ 5.2 Analysis of Data Organization Paradigms
We compare various data organization paradigms within RAG, implementing Semantic Chunking Kamradt (2024) and Proposition Chen et al. (2023) in addition to recursive chunking Splitter (2023) (chunk-based) and entire document (doc-based) approaches (details in Appendix A). Table 5 shows that our Thread paradigm outperforms the other paradigms across all metrics. Although Semantic and Proposition methods use LLMs to merge semantically similar sentences, they fail to adequately address logical connections, resulting in poor performance. Additionally, our system processes the smallest token length of retrieval units yet achieves the highest performance, underscoring our approach’s efficiency and effectiveness.
@@
Table 5: Analysis of data organization paradigms. Threadw/o. represents Thread without LU selector.
5.3 Document Formats in LU Extraction
As mentioned in §4.2, we tested our LU extraction method in accommodating varying document structures. We generated different document formats (detailed in Appendix E.2), including structured markdown, hierarchical guidelines, tabular checklists, and narrative documents. Table 6 shows that our Thread paradigm effectively organizes these formats, consistently outperforming the chunk-based paradigm across all metrics. Our Thread improves Ele. Acc by up to 9.61%, Op. F1 by up to 8.43%, and Step SR by up to 8.51%. Notably, our Thread achieves the highest performance with structured documents, as this format helps construct a higher-quality knowledge base.
@@ 6 Conclusion
In this paper, we address the overlooked category of how-to questions within existing QA systems. We propose Thread, a novel data organization paradigm to capture logic connections with documents. Thread is compatible and integrated to current RAG-based QA system. Experimental results demonstrate Thread’s effectiveness in handling how-to questions, surpassing other data organization paradigms in performance.
Limitations
This work choose three scenarios to design how-to questions, and evaluate its effectiveness of Thread on handling both linear how-to questions and dynamic how-to questions. However, the proposed paradigm still has some limitations for future directions. On the one hand, since our logic knowledge base can coexist with the original chunk knowledge base, we do not extend our method to test factoid questions, like multi-hop questions, long-form questions and so on. On the other hand, the experiments primarily rely on OpenAI’s backbone models. Future studies should include evaluations using other LLMs like LLaMA and retrievers such as Contriver to better validate the effectiveness of our paradigm. Additionally, while extracting logic units involves initial costs in calling LLMs, our approach is a one-time procedure. Once the knowledge base is constructed, it becomes advantageous for industrial applications, particularly in terms of subsequent updates and maintenance.
Ethics Statement
All the experiments are conducted on existing dataset or resources that collected by ourselves or from internal sources. We keep fair and honest in our analysis of experimental results. As the scenario of IcM need human evaluation, we ensure that there is no bias in selecting OCEs and ensure the randomness of the sampling.
Appendix A Details about Current Data Organization Paradigms
@@
In our experiments, we set the temperature of LLMs to 0 and top_p to 1 for results reproduction.
Appendix B Examples of Each Dataset
Table B shows the examples of each dataset.
@@ Appendix B Examples of Each Dataset
### Step 1: Check Pull Task Execution From the Cluster The direct impact of connection failure is pull task execution will not work. If Service A can continue to pull from Service B, then the incident can be dismissed as false alarm, the feature owner can investigate further to see why Echo fails. This can be visualized by pull task count over time in the last 8 hours in the following query: ‘‘‘kusto <Code Block> ‘‘‘ Disregard the last data point, if the data point is always above zero, then consider the alert as false alarm. If the chart sometimes drops to zero one hour ago and the number is low in general (for instance less than 20), it means the customer traffic in the cluster is low. In this case, observe for a longer period of time. If the data point is zero consistently in the past 30 minutes, then it is a real problem, and please Check if Other Clusters In the Region are Impacted. Otherwise, continue to observe since Service A is pulling Service B just fine.
@@
Appendix C Example of Logic Unit
Table C shows an example to demonstrate the reformulated process of the document and its corresponding logic unit.
Trouble Shooting Guide: How to Investigate Service A-To-Service B Connection?
Original ### Step 0: Determine the Region and Cluster Name The region and cluster name can be found in the incident title. ### Step 1: Check Pull Task Execution From the Cluster The direct impact of connection failure is pull task execution will not work. If Service A can continue to pull from Service B, then the incident can be dismissed as false alarm, the feature owner can investigate further to see why Echo fails. This can be visualized by pull task count over time in the last 8 hours in the following query: *** Disregard the last data point, if the data point is always above zero, then consider the alert as false alarm. If the chart sometimes drops to zero one hour ago and the number is low in general (for instance less than 20), it means the customer traffic in the cluster is low. In this case, observe for a longer period of time. If the data point is zero consistently in the past 30 minutes, then it is a real problem, and please Check if Other Clusters In the Region are Impacted. Otherwise, continue to observe since Service A is pulling Service B just fine.
Reformulated ## 1.Check Pull Task Execution From the Cluster. ### Prerequisite The region and cluster name can be found in the incident title. ### Header Check Pull Task Execution From the Cluster ### Body Run the following query to check pull task execution from the cluster (please use the cluster name from the previous step) *** ### Linker - If the data point is always above zero, then consider the alert as false alarm.[MITIGATE] - If the chart sometimes drops to zero one hour ago and the number is low in general, it means the customer traffic in the cluster is low. In this case, observe for a longer period of time.[MITIGATE] - If the data point is zero consistently in the past 30 minutes, then it is a real problem, and please Check if Other Clusters In the Region are Impacted.[CONTINUE] - Otherwise, continue to observe since Service A is pulling Service B just fine.[MITIGATE]
Logic Unit { "#type#": "step", "#meta data#": { "#title#": "How to Investigate Service A-To-Service B Connection", "#id#: "", "#date#": "" }, "#prerequisite#": "The region and cluster name are given.", "#header#": "Check Pull Task Execution From the Cluster.", "#body#": "Run the following query to check pull task execution from the cluster (please use the cluster name from the previous step):***", "#linker#": "If the data point is always above zero, then consider the alert as false alarm.[MITIGATE] If the chart sometimes drops to zero one hour ago and the number is low in general, it means the customer traffic in the cluster is low. In this case, observe for a longer period of time.[MITIGATE] If the data point is zero consistently in the past 30 minutes, then it is a real problem, and please Check if Other Clusters In the Region are Impacted.[CONTINUE] Otherwise, continue to observe since Service A is pulling Service B just fine.[MITIGATE]", "#default_parameters#": { "<TIME>": "", "<CLUSTER NAME>": "" } } ...
Table 8: An example of reformulated TSG and its corresponding Logic Unit using Thread.
Appendix D Details about Incident Mitigation
We take the scenario of incident mitigation to show the instructions of how to construct our knowledge base, including reformulation, code template extraction, and selection.
D.1 Document Reformulation Instruction
Table 9 shows the instructions we use to reformulate the document.
[System] You are a helpful troubleshooting guide assistant who helps the user formulate the manual unstructured troubleshooting guide <TSG> into a structured one. The <TSG> is in markdown format, with the first level header describing the incident or problem, and the following second level header providing information related to the incident or problem. Each second-level subsection can be categorized into the following types: Terminology, FAQ, STEP, and Appendix. Your reformulation should strictly comply with the following definition: - Terminology: firstly, it should be the relationship or connection between terminology about the incident, if not, it can be the explanation or concept of the incident. Sometimes it should be extracted and summarized by yourself. - FAQ: frequently asked questions that help to understand the incident. - STEP: the processes to resolve the incident, and you should make sure its completeness. Usually, steps have causal inner connection, the former step will trigger the next step. - Appendix: the supplement of the incident that is not important or labeled by TSG, usually providing additional resources, data, links and so on. 1. You need to identify each second-level subsection, including third-level subsection if needed, analyze its content or purpose, and categorize it accordingly. For those belonging to Step, you should capture the inner connections, such as Causality or Temporal relations, and present them in the correct order. 2. Your returned formulated TSG should be in JSON format. Make sure that the keys originate from these categories: Terminology, FAQ, STEP ad Appendix. Each value should be a list of dictionaries. The keys for them are "prerequisite", "header", "body", and "linker". All values within the lists need to align with the original context, with truthful meaning and necessary **code block**. 3. Importantly, the "linker" is used to imply the dual role of providing the action’s result and connecting to the next step using the "if-then" sentence format. You should formulate each step’s linker to be "If any results are obtained by executing the corresponding action in the previous step, then **the true intent of the following step** provided here". Implicit linkers like "proceed to the next step." or "then the intent of the following step should be taken into consideration." should be avoided. 4. For each "if" condition at every step in the STEP, it is necessary to add a special token behind the "then" condition within the "linker". The options for these tokens are "[CONTINUE]", "[CROSS]", and "[MITIGATE]". - The token "[CONTINUE]" indicates that the actions corresponding to this "if" condition are part of the continuum within the same TSG’s STEPs. - The token "[CROSS]" signifies that the subsequent actions require a transition to a different set of steps that are external to the current TSG’s STEPs. - The token "[MITIGATE]" implies that the actions following the "if" condition convey that the incident is mitigated, or necessitate communication with on-call engineers or teams. The use of this special token is instrumental in verifying the completeness and structural integrity of the STEP section. <TWO EXAMPLES HERE> [User] Here is the <TSG> you need to formulate: {TSG}
@@ D.1 Document Reformulation Instruction
Table 9: Instruction that formulates the original unstructured troubleshooting guide into structured one.
D.2 Code Template Instruction
Table 10 shows the instruction we use to extract the code template and default parameters from the original code.
[System] You are a helpful assistant that extracts the code template and the default parameters from the provided code instance in <CODE>. <CODE> is a code block that contains several parameters. You should replace those parameters with placeholders and output the code template with placeholders and default parameters. <ONE EXAMPLE HERE> Your response should be in the JSON format as below: { "#CODE_TEMPLATE#": where you replace the parameters in <CODE> with placeholders, "#DEFAULT_PARAMETERS#": where you keep the parameters in <CODE> as default values. } [User] Here is the <CODE> you need to extract: {CODE}

### v2

@@ 4.1 Scenarios and Datasets
We evaluate Thread on two open-domain scenarios: Web Navigation (Mind2Web (Deng et al., 2023a)) and Wikipedia Instructions (WikiHow (Koupaee & Wang, 2018)), and one industrial setting: Incident Mitigation (IcM (Shetty et al., 2022; An et al., 2024)). We provide one example of each dataset in Appendix B to better demonstrate linear and dynamics how-to questions.
Web Navigation. Mind2Web (Deng et al., 2023a) is a dataset designed for web agents to perform complex tasks on real-world websites based on language instructions. Each task involves a ‘dynamic how-to question’, with multiple possible outcomes depending on the state of the executed actions.
Wikipedia Instructions. WikiHow [^4] https://www.wikihow.com is a platform containing numerous articles that offer step-by-step guidance on various procedural tasks. Each article is typically titled ‘How to’ and includes a brief task description, followed by a linear, fixed sequence of steps.
Incident Mitigation. Incident Mitigation (Shetty et al., 2022; An et al., 2024; Jiang et al., 2024) is essential for operating large-scale cloud services, where engineers use Troubleshooting Guides (TSGs) to address incidents. Each step in incident mitigation can yield different outcomes depending on the state, making it suitable for testing Thread on ‘dynamic how-to questions’. Unlike the other open-domain datasets, we perform a human evaluation involving twenty on-call engineers (OCEs) responsible for incident mitigation. We collect five incidents, classified into two simple and three hard ones based on their mitigation steps in history. Each OCE is tasked with mitigating all five incidents, using randomly ordered baselines per incident to avoid familiarity bias. The RAG system initiates automated mitigation for each incident; if it encounters a failure at any step, an OCE intervenes to address the issue before the system resumes automated procedures [^5] Varying across new-hire and experienced OCEs. Mitigating one incident costs each OCE around 30 minutes to 1 hour manually, and we collect five incidents to ensure consistency, engagement, and ethical treatment. Note that one OCE’s data was contaminated during the experiment, so we removed that OCE’s data..
@@ 4.2 Documents for Retrieval
Table 1: The statistics of datasets, including the number of documents, derived chunks, LUs, etc.
As the Mind2Web dataset lacks relevant documents, we create retrieval documents tailored to the RAG system. Assuming that each website has help docs applicable across different tasks, we use the "Cross-Task" test set, selecting examples from the training set to craft informative documents for each website. And we follow Wang et al. (2023) to brainstorm different formats of documents. More details about collecting documents are shown in Appendix C.2.
For the WikiHow dataset, we utilize publicly available Windows Office Support Docs [^6] https://github.com/MicrosoftDocs/OfficeDocs-Support as retrieval documents. We select around 100 tasks from WikiHow tagged with Microsoft products like Word, PowerPoint, and Teams, each containing 10 to 40 steps related to Windows operations.
For the IcM dataset, we collect 56 TSGs from an enterprise-level engineering team responsible for a large-scale cloud platform. The selected incidents in the IcM dataset can be resolved using the knowledge provided in these TSGs. We list the statistics of our datasets in Table 1.
@@ 4.3 Baselines
In the Mind2Web dataset, previous work has not treated it as how-to questions. State-of-the-art methods like SYNASE (Zheng et al., 2023) and MINDACT (Deng et al., 2023a) either use In-Context Learning (ICL), providing few-shot demonstrations, or Supervised Learning (SL) to finetune a model on the training set. To ensure a fair comparison with our LLM endpoints [^7] We use GPT-3.5 and GPT-4 with version 1106-preview., we re-implement the MINDACT experiments with the same demonstrations and included chat history as extra context. In our paper, we treat the Mind2Web dataset as dynamic how-to questions and use our RAG system to solve these tasks. For comparison, we use doc-based (providing the entire document) and chunk-based data organization paradigms as baselines against Thread. This RAG system with different data organization paradigms also serves as baselines for the WikiHow and IcM datasets. More details about experiments can be found in Appendix LABEL:appendix_expD.
@@ 4.4 Evaluation Metrics
For the Mind2Web dataset, each task is treated as a multi-choice question. We adapt the evaluation metrics used in (Deng et al., 2023a), which include: Element Accuracy (Ele. Acc) to compare the chosen HTML element with all provided elements; Operation F1 (Op. F1) to calculate the token-level F1 score for predicted operations such as “CLICK”, “TYPE IN”, etc.; Step Success Rate (Step SR), where a step is successful if both the selected element and predicted operation are accurate; and Success Rate (SR), where a task is successful only if all steps are successful.
For the WikiHow dataset, which contains ground truth steps, we leverage LLMs to extract “Action Items” from each ground truth step and generated step, and we use the following metrics: Precision ($P=\frac{\#matched\_items}{\#total\_generated\_items}$); Recall ($R=\frac{\#matched\_items}{\#total\_groundtruth\_items}$); F1; and Success Rate (SR) to assess if the generated steps can successfully complete the task, using LLMs to evaluate (Appendix C.3 shows the evaluation prompt).
@@ 5 Experimental Results
5.1 Main Results
Web Navigation. The overall performance on Mind2Web is shown in Table 2, where we compare our method with baselines and both doc-based and chunk-based RAG methods. We observe that providing informative documents (regardless of the data organization paradigm) within RAG methods significantly improves performance. RAG methods outperform ICL and SL methods. By incorporating external documents, the doc-based RAG method achieves performance comparable to the best results of MINDACT. Notably, our Thread paradigm is the best among RAG methods, showing improvements of 4.06% in Ele. Acc, 3.49% in Step SR and 3.57% in SR. Note that MINDACT-SL gets the highest Op. F1 due to label distribution imbalance [^8] If the model predicts all operations as ”CLICK,” the Op. F1 would reach 79.90%., leading the model to favor generating the most frequent operations.
@@ 5.1 Main Results
Table 2: Experiment results on Mind2Web. ‘*’ represents taking results from the original paper.
Incident Mitigation
Paradigm Category SR Step SR P.F. Step SR HI Turns
Simple 40.51 60.90 60.90 30.10 3.14
Chunk Hard 28.95 53.16 43.05 46.84 6.84
Simple 43.86 63.90 63.90 36.09 2.98
Doc Hard 31.58 57.89 42.11 42.11 6.53
Simple 77.19 88.72 84.21 11.28 2.56
Thread Hard 52.63 84.21 68.95 15.79 5.74
Table 3: Experiment results on Incident Mitigation. We divide incidents into two groups: simple or hard, and compare the performance separately.
Incident Mitigation. Table 3 illustrates the advantages of Thread over other paradigms when addressing complex dynamic how-to questions. Both chunk-based and doc-based RAG methods exhibit significant limitations in incident mitigation, as their SR and P.F. Step SR are much lower due to their inability to connect subsequent steps based on the current step. In contrast, our Thread achieves the best score across all metrics, particularly achieving a significant increase in SR from 21.02% to 33.33%. More importantly, the highest P.F. Step SR highlights Thread’s ability to dynamically connect to subsequent steps based on user feedback, thereby reducing the need for human intervention. Consequently, Thread not only achieves the highest Step SR but also requires the fewest interaction turns to effectively mitigate both simple and complex incidents.
@@
Table 4: Experiment results on WikiHow with different interaction manners and paradigms.
Wikipedia Instructions. Table 4 presents the performance of various data organization paradigms on WikiHow tasks, experimenting with both single-turn and multi-turn interactions. In the single-turn setting, where the RAG system needs to generate the entire plan in a single interaction, our Thread outperforms the doc-based method with 5.15%. This indicates that Thread delivers information with less redundancy while achieving a higher SR score. Additionally, when comparing Thread to the chunk-based method, we observe a significantly lower success rate (SR) of 19.59%. This result highlights the challenge of retrieving all relevant chunks without considering the logic flow inside documents.
In the multi-turn setting, the RAG system performs iterative retrieval, showing superior performance compared to single-turn. This indicates the advantage of a step-by-step approach in handling how-to questions. We can see that the SR improves significantly from 58.76% to 68.04% in the doc-based paradigm and from 63.91% to 72.16% in Thread. Similar to the single-turn setting, the chunk-based method splits documents while disrupting internal logic, resulting in a low SR of 20.62%. Our Thread excels across all metrics, achieving the highest SR of 72.16%, highlighting its effectiveness in maintaining and modeling the connections between steps.
5.2 Ablation on RAG System Settings
This section presents an ablation study on the key settings of our RAG system. While retriever and generator variants have been explored in other research (Gao et al., 2023), we use text-embedding-ada-002 (OpenAI, 2022) for the retriever and GPT-4 as the LLM-based generator. The ablation study is performed on Mind2Web dataset.
Multi-turn Interaction. As shown in Table 4, the multi-turn setting is better than the single-turn setting in answering how-to questions. We use multi-turn setting for the other scenarios in this paper.
Chat History. The chat history helps the system to check previous actions and results, helping the system make better decisions. Without chat history, the performance drops as shown in Table 5. We include chat history for all RAG-based methods in our experiments.
@@ 5.2 Ablation on RAG System Settings
Table 5: Ablation study of integrating chat history and retrieval unit selection on Mind2web.
Retrieval Units Selector. As aforementioned in §3.4 and Figure 3, the Selector picks the most relevant retrieval units from the top-K retrieved units. Table 5 shows the ablation of the retrieval unit selector. Comparing chunk selection to LU selection, the performance of Ele. Acc drops by 4.29% and Op.F1 by 3.38%. Conversely, the selector improves all metrics with applied to Thread. Unlike LU selection, which filters out irrelevant LUs based on prerequisites, chunk selection disregards inter-chunk connections and may exclude relevant chunks. Therefore, we activate the selector only for our Thread.
@@ 5.3 Analysis of Data Organization Paradigms
We compare various data organization paradigms within RAG, implementing Semantic Chunking (Kamradt, 2024) and Proposition (Chen et al., 2023) in addition to recursive chunking (Splitter, 2023) (chunk-based) and entire document (doc-based) approaches (details in Appendix A). Table 7 shows that our Thread paradigm outperforms the other paradigms across all metrics. Thread w/o. represents without LU selector. Although Semantic and Proposition methods use LLMs to merge semantically similar sentences, they fail to adequately address logical connections, resulting in poor performance. Additionally, our system processes the smallest token length of retrieval units yet achieves the highest performance, underscoring our approach’s efficiency and effectiveness.
Figure 4: Analysis of using different LLMs on WikiHow (GPT-4 v.s. LLaMA3-70B).
@@
Table 6: Analysis of data organization paradigms.
@@
5.4 Document Formats in LU Extraction
As mentioned in §4.2, we test our LU extraction method in accommodating varying document structures. We generate different document formats (detailed in Appendix C.2), including structured markdown, hierarchical guidelines, tabular checklists, and narrative documents. Table 7 shows that our Thread paradigm effectively organizes these formats, consistently outperforming the chunk-based paradigm across all metrics. Our Thread improves Ele. Acc by up to 9.61%, Op. F1 by up to 8.43%, and Step SR by up to 8.51%. Especially, our Thread achieves the highest performance with structured documents, as this format helps construct a higher-quality knowledge base.
5.5 Different LLMs as backbone
We further conduct experiments on LLaMA3-70B, as shown in Figure 4. Although LLaMA3-70B is less powerful than GPT-4, it still demonstrates competitive performance with the help of Thread. Results indicate that LLaMA3 struggles with WikiHow questions applying chunk-based or doc-based paradigms. However, with the integration of Thread, LLaMA3 not only achieves much better performance in SR, but also narrows the gap with GPT-4. Specifically, LLaMA3 achieves SR and F1 scores of 60.82% and 72.84%, respectively, compared to GPT-4’s scores of 72.16% and 80.74%. This indicates that while a performance gap remains, particularly with the chunk-based and doc-based paradigms, our Thread considerably reduces this disparity between GPT-4 and LLaMA3. These findings highlight the value of our proposed paradigm in enhancing the performance of different LLMs, showcasing its generalizability, robustness, and efficiency in handling how-to questions.
@@ 6 Conclusion
In this paper, we address the overlooked category of handling how-to questions in QA systems by proposing Thread, a novel data organization paradigm that captures logical connections within documents. By introducing a new knowledge granularity called ’logic unit,’ Thread restructures documents into interconnected logic units that are compatible with RAG methods. Extensive experiments show that Thread significantly outperforms existing paradigms, improving performance while reducing the knowledge base size and minimizing the information needed for generation.
Appendix A Current Data Organization Paradigm
@@
In all our experiments, we set the temperature of LLMs to 0 and top_p to 1 for results reproduction.
Appendix B Example of Dataset
We list one example of each dataset in Table B, and an example of the construction process of Thread including the original document, reformulated document, and its corresponding logic unit in Table LABEL:tab:example_lu.
@@ Appendix B Example of Dataset
How to Investigate Service A-To-Service B Connection? ### Step 1: Check Pull Task Execution From the Cluster The direct impact of connection failure is pull task execution will not work. If Service A can continue to pull from Service B, then the incident can be dismissed as false alarm, the feature owner can investigate further to see why Echo fails. This can be visualized by pull task count over time in the last 8 hours in the following query: ‘‘‘kusto <Code Block> ‘‘‘ Disregard the last data point, if the data point is always above zero, then consider the alert as false alarm. If the chart sometimes drops to zero one hour ago and the number is low in general (for instance less than 20), it means the customer traffic in the cluster is low. In this case, observe for a longer period of time. If the data point is zero consistently in the past 30 minutes, then it is a real problem, and please Check if Other Clusters In the Region are Impacted. Otherwise, continue to observe since Service A is pulling Service B just fine.
@@
- Terminology: firstly, it should be the relationship or connection between terminology about the incident, if not, it can be the explanation or concept of the incident. Sometimes it should be extracted and summarized by yourself. - FAQ: frequently asked questions that help to understand the incident. - STEP: the processes to resolve the incident, and you should make sure its completeness. Usually, steps have causal inner connection, the former step will trigger the next step. - Appendix: the supplement of the incident that is not important or labeled by TSG, usually providing additional resources, data, links and so on.
1. You need to identify each second-level subsection, including third-level subsection if needed, analyze its content or purpose, and categorize it accordingly. For those belonging to Step, you should capture the inner connections, such as Causality or Temporal relations, and present them in the correct order.
2. Your returned formulated TSG should be in JSON format. Make sure that the keys originate from these categories: Terminology, FAQ, STEP ad Appendix. Each value should be a list of dictionaries. The keys for them are "prerequisite", "header", "body", and "linker". All values within the lists need to align with the original context, with truthful meaning and necessary **code block**.
3. Importantly, the "linker" is used to imply the dual role of providing the action’s result and connecting to the next step using the "if-then" sentence format. You should formulate each step’s linker to be "If any results are obtained by executing the corresponding action in the previous step, then **the true intent of the following step** provided here". Implicit linkers like "proceed to the next step." or "then the intent of the following step should be taken into consideration." should be avoided.
4. For each "if" condition at every step in the STEP, it is necessary to add a special token behind the "then" condition within the "linker". The options for these tokens are "[CONTINUE]", "[CROSS]", and "[MITIGATE]". - The token "[CONTINUE]" indicates that the actions corresponding to this "if" condition are part of the continuum within the same TSG’s STEPs. - The token "[CROSS]" signifies that the subsequent actions require a transition to a different set of steps that are external to the current TSG’s STEPs. - The token "[MITIGATE]" implies that the actions following the "if" condition convey that the incident is mitigated, or necessitate communication with on-call engineers or teams.
The use of this special token is instrumental in verifying the completeness and structural integrity of the STEP section.
<TWO EXAMPLES HERE>
[User]
Here is the <TSG> you need to formulate: {TSG}
@@
Instruction that formulates the original unstructured troubleshooting guide into structured one.
<ONE EXAMPLE HERE>
Your response should be in the JSON format as below:
"#CODE_TEMPLATE#": where you replace the parameters in <CODE> with placeholders,
"#DEFAULT_PARAMETERS#": where you keep the parameters in <CODE> as default values.
[User]
Here is the <CODE> you need to extract: {CODE}